a certified public accountant (cpa) randomly selects 4 accounts for a group of 12 careful examination. The company being audited knows that 7 of the accounts contain - errors. what is the probability that all 3 accounts selected are error-free


1
Expert's answer
2022-02-16T09:06:53-0500

Solution:

Assume total no. of accounts the company has 14. (It is missing in question).

Total no. of selected accounts = 3

No. of accounts contain - errors = 7

No. of accounts errors free = 14 - 7 = 7

Required probability "=\dfrac{^7C_3}{^{14}C_3}=\dfrac{35}{364}=\dfrac{5}{52}"